OpenStudy (anonymous):

Jim has an annual salary of $96,000. His monthly expenses include a $2,500 mortgage payment, a $250 lease payment, $500 in minimum credit card payments, and a $425 payment on his speed boat. He also receives $1,200 in interest from his savings and other accounts each month. Calculate Jim’s DTI (debt-to-income) ratio.

OpenStudy (dumbcow):

total debt : 2500 + 250 +500+425 = 3675 total income : 8000 + 1200 = 9200 debt/income = 3675/9200